The Village BBS

Florida Arcade and Pinball Collectors => Technical Questions and Help => Topic started by: WannaTheater on January 14, 2013, 02:36:53 PM

Title: No fear opto issue
Post by: WannaTheater on January 14, 2013, 02:36:53 PM
I am pulling my hair out - There is a subway under the No Fear playfield that had 4 sets of optos.  In the test menu, the "Popper 2" opto set is showing as X'ed (i.e. something is either blocking the path of the IR, or the sender and receiver are out of alignment.  I have swapped 3 of the sets of optos around, hoping that the problem could be isolated to that pair of optos.  It is not.  It appears that all 4 sets of optos are fine - but the pin still thinks that the set in "Popper 2" is bad.

I have traced all wires from the pair of optos (send/receive) in question to the under-playfield opto processing board, and they are fine.  I have also reflowed the header solder pins on the opto processing board.

But still the CPU sees the opto pair as non-operational.  Any ideas?
Title: Re: No fear opto issue
Post by: WannaTheater on January 14, 2013, 03:23:20 PM
Actually, now 2 sets are bad - Popper 1 and Popper 2.

Title: Re: No fear opto issue
Post by: WannaTheater on January 14, 2013, 10:02:04 PM
It was 2 bad IR transmitters.  I was able to purchase the IR LED emitters from Radio Shack which fixed the problem.

Title: Re: No fear opto issue
Post by: k7 on January 15, 2013, 12:21:14 AM
nice. radio shack actually comes through with a local part. :)